OpenStudy (anonymous):

Find all the zeros. x^3+x^2=7x+7. Show all work.

OpenStudy (jhannybean):

@WesternBurger ?

OpenStudy (anonymous):

Yeah?

OpenStudy (jhannybean):

Alright :) When finding zeroes, are you looking for real and complex,or just the real?

OpenStudy (anonymous):

Uhmm both? And hey I'm gonna close this question but I'm still gonna respond and everything. I just need to post a few more questions.

OpenStudy (jhannybean):

Hm,how about finishing this one,then opening up a new one?Just a suggestion...

OpenStudy (anonymous):

I know I usually do but I'm in a time crunch. It's almost 5:30 am I've been up since 4 and I need to start getting ready for school soon.

OpenStudy (jhannybean):

I see.This one won't take very long :)

OpenStudy (jhannybean):

Let's work on this. Move everything over to the left side. what is your equation?

OpenStudy (anonymous):

It would be... x^3+x^2-7x-7=0

OpenStudy (jhannybean):

Awesome,now group together two terms at a time. \[(x^3 +x^2) +(-7x-7)=0\]

OpenStudy (jhannybean):

Factor out your LCD.

OpenStudy (anonymous):

x^2(x+1) -7 (x+1) (x^2-7) (x+1) So the zeros would be -1 and + or - square root of 7?

OpenStudy (jhannybean):

Actually,you got themboth right.

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

Would it be \[\pm \sqrt{7}\] Maybe I just worded it wrong lol

OpenStudy (jhannybean):

I read it wrong at first.

OpenStudy (anonymous):

Oh ok, thanks so much! :)

OpenStudy (jhannybean):

No problem :)